# Topologically protected negative entanglement

### v3

@@ I.1.2 Effect of bulk states on entanglement scaling due to topological EPs
In the preceding subsection II.B, we demonstrated that near $k=0$, the dominant element of $P(k)$ is located in the upper right corner, where $P_{1,2L_{y}}(k)\sim 1/r^{L_{y}}$ with $r=\sqrt{\big{|}\frac{(b_{0}-\cos k)^{B}}{a_{0}}\big{|}}<1$. For $b_{0}>1$, this can still be the case with appropriate choice of $a_{0}$. We have
$\bra{x_{1}}\bar{P}_{1,2L_{y}}\ket{x_{2}}=\frac{1}{L}\sum_{k}e^{ik(x_{1}-x_{2})}P_{1,2L_{y}}(k)\quad\sim\frac{1}{L}\sum_{k}e^{ik(x_{1}-x_{2})}a_{0}^{L_{y}/2}\frac{1}{(b_{0}-\cos k)^{BL_{y}/2}},$ (S47)
@@
When $\color[rgb]{0,0,1}L\ll L_{y}$, (for simplicity, we select $L_{y}$ to be an even number to ensure $BL_{y}/2$ is an integer), the following approximation holds
$(1+x)^{BL_{y}}=\sum_{m}C_{BL_{y}/2}^{m}x^{m}\approx C_{BL_{y}/2}^{M}x^{M}$ (S49)
where $m=M$ represents the maximum term in the summation. As verified in Fig. S5(a), $M$ exhibits a linear relation with $L_{y}$, expressed as $M\approx\frac{1}{2}(\gamma L_{y}+\zeta)$ (where $\gamma,\zeta$ are coefficients independent of $L_{y}$). Consequently, we find
$(1+x)^{BL_{y}}\approx C_{BL_{y}/2}^{M}x^{M}\sim C_{BL_{y}/2}^{(\gamma L_{y}+\zeta)/2}\left(\frac{\pi^{2}}{2(b_{0}-1)L^{2}}\right)^{(\gamma L_{y}+\zeta)/2}\sim L^{-(\gamma L_{y}+\zeta)}.$ (S50)
By substituting into Eq. (S48) and (S47) and considering only the contribution from $k_{1}=\pi/L$, we have
$\bra{x_{1}}\bar{P}_{1,2L_{y}}\ket{x_{2}}\sim\frac{1}{L}e^{ik_{0}(x_{1}-x_{2})}\frac{a_{0}^{L_{y}/2}}{(b_{0}-1)^{BL_{y}/2}(1+x)^{BL_{y}/2}}\sim L^{\gamma L_{y}+\zeta-1}.$ (S51)
As we have demonstrated in the previous subsection B, since the elements of $\bar{P}_{1,2L_{y}}$ are proportional to $L^{\gamma L_{y}+\zeta-1}$, we can deduce:
$\text{Tr}(\bar{P}^{2})$$\approx$ $2\text{Tr}(\bar{P}_{1,2L_{y}}\bar{P}_{2L_{y},1})\sim L^{\gamma L_{y}+\zeta-1}\Longrightarrow\quad p_{i}^{1st}\sim L^{\frac{\gamma L_{y}+\zeta-1}{2}}$
$S_{A}$ $\sim$ $-4\times\left(\frac{\gamma L_{y}+\zeta-1}{2}\right)\log L\sim-(\alpha L_{y}+\beta)\log L,$ (S52)
where we have redefined $2\gamma$ as $\alpha$ and $2(\zeta-1)$ as $\beta$. In the $L\ll L_{y}$ regime, higher orders of $p_{i}$ are negligible, and we only need to consider the first order. Therefore, we conclude: for $L\ll L_{y}$ , $S_{A}\sim-\log L$ (Fig. S5(b)), with the gradient depending linearly on $L_{y}$ (see Fig. S5(c)). Even though our derivation had assumed $L\ll L_{y}$, from the numerics in Fig. 2(e) of the main text, we see that this trend still mostly holds as long as $L<L_{y}$.
Figure S5: Negative entanglement scaling for the gapped EP case ($b_{0}>1$). (a) As numerically plotted for $L=9,b_{0}=1.2$, the approximate linear scaling of the scaling exponent $M$ of the dominant $\bar{P}$ eigenvalue with $L_{y}$, as described by Eq. (S49). (b) The negatively linear relationship of the entanglement entropy $S_{A}$ with $\log L$ for different $L_{y}$ ranging from 22 to 32, with $L\in[22,32]$ and parameters $t=0.8,a_{0}=1,b_{0}=1.2,B=1$. As predicted in Eq. (S52), the gradient indeed increases with $L_{y}$. (c) Using the same parameters as in (b), approximate linear dependence of the gradient of $S_{A}$ on $L_{y}$. (d) For large $L=60$ and parameters $t=0.8,B=1,b_{0}=1.2,a_{0}=2$, $S_{min}$ indeed scales linearly with $L_{y}$, as predicted in Eq. (S55).
When $\color[rgb]{0,0,1}L\gg L_{y}$, Eq. (S48) can be approximated as:
$(b_{0}-\cos k_{0})^{BL_{y}/2}$ $=$ $\left(b_{0}-1+\frac{1}{2!}\left(\frac{\pi}{L}\right)^{2}-\frac{1}{4!}\left(\frac{\pi}{L}\right)^{4}+\cdots\right)^{BL_{y}/2}$ (S53)
@@
Thus, Eq. (S47), which represents the elements of the largest matrix block $\bar{P}_{1,2L_{y}}$, can be expressed as
$\bra{x_{1}}\bar{P}_{1,2L_{y}}\ket{x_{2}}\sim\left(\frac{a_{0}}{(b_{0}-1)^{B}}\right)^{L_{y}/2}.$ (S54)
Following a similar derivation as Eq. (S52), we now have
@@
$S_{min}$ $\sim$ $-4\times\frac{L_{y}}{4}\log\left(\frac{a_{0}}{(b_{0}-1)^{B}}\right)=-L_{y}\log\left(\frac{a_{0}}{(b_{0}-1)^{B}}\right).$ (S55)
@@
This linear dependence on $L_{y}$ is verified in Fig. S5(d). Empirically, we also see from Fig. 2(e) of the main text that this saturation generally sets in as long as $L$ exceeds $L_{y}$.
III III. Further properties of the negative entanglement
III.1 A. Negativity of the Rényi entropy
Below, we show that when the entanglement entropy is negative, so is the Rényi entropy. For $|p_{i}|\gg 1$, we observe that $p_{i}^{n}+(1-p_{i})^{n}\propto|p_{i}|^{n}\gg 1$ for even-order $n$, and $p_{i}^{n}+(1-p_{i})^{n}\propto|p_{i}|^{n-1}\gg 1$ when $n$ is odd. Substituting this $p_{i}$ into Eq. (3) makes it is evident that the $n$th-order Rényi entropy will be negative. Similarly, for the von Neumann entropy as detailed in Eq. (4), the contribution of $p_{i}$ to the real part of $S_{A}$, when $p_{i}\gg 1$, is given by
$\text{Re}[S_{A}(p_{i})]$ $=$ $-p_{i}\log(p_{i})+(p_{i}-1)\log(p_{i}-1)<0,$ (S56)
and for $p_{i}\ll 0$, the contribution is
$\text{Re}[S_{A}(p_{i})]$ $=$ $|p_{i}|\log(|p_{i}|)-(|p_{i}|+1)\log(|p_{i}|+1)<0.$ (S57)
Thus, eigenvalues with $|p_{i}|\gg 1$ significantly contribute to a negative value for both the von Neumann and the nth-order ($n\geq 2$) Rényi entropies.
@@
III.2 B. PT-symmetry-protected real entanglement entropy $S_{A}$ in the flat band model
@@
In this section, we will demonstrate that the entanglement entropy $S_{A}$ is rigorously real for PT-symmetry-protected models, including our topological flat band model. This can be attributed to the fact that the eigenvalues, $p_{i}$, of $\bar{P}$ consistently appear as complex conjugate pairs.
@@
First, in the $H_{\text{y-OBC}}(k)$ model of Eq. (S25), the parameters $a_{0},b_{0},t$ are always real numbers, with $b_{0}\geq 1$. As previously demonstrated, this model can be viewed as a non-Hermitian SSH model that is protected by PT-symmetry, and it features purely real eigenvalues and eigenvectors, $|\psi_{n}^{R}(k)\rangle$ and $\langle\psi_{n}^{L}(k)|$. Therefore, the elements of the projector matrix $P_{\alpha,\beta}(k)=\sum_{n\in\text{occ}}\langle\alpha|\psi_{n}^{R}(k)\rangle\langle\psi_{n}^{L}(k)|\beta\rangle$ are all real, with $\alpha,\beta$ labeling the band indices. Upon expanding these elements into real space, and given that $P_{\alpha,\beta}(k)$ are real, we will obtain
$\langle x_{1}|P_{\alpha,\beta}|x_{2}\rangle=\sum_{k}P_{\alpha,\beta}(k)e^{ik(x_{1}-x_{2})}=\Big{(}\langle x_{2}|P_{\alpha,\beta}|x_{1}\rangle\Big{)}^{*}.$ (S58)
@@
$V\bar{P}V^{-1}=\text{conj}(\bar{P}),\text{ with }V=\begin{pmatrix}0&0&0&\cdots&\mathbb{I}\\ 0&0&\cdots&\mathbb{I}&0\\ 0&\cdots&\mathbb{I}&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \mathbb{I}&0&0&\cdots&0\end{pmatrix},$ (S59)
@@
indicating that $\bar{P}$ possesses time-reversal symmetry and its spectrum must consists of complex conjugate pairs $(p,p^{*})$.
In general, for such conjugate pairs of $\bar{P}$ eigenvalues $(p,p^{*})$ as $(p_{a}e^{i\theta_{a}},p_{a}e^{-i\theta_{a}})$, where $p_{a},\theta_{a}$ are real numbers, their contribution to the entanglement entropy can be expressed as
$S_{\text{pair}}$ $=$ $-p_{a}e^{i\theta_{a}}\log(p_{a}e^{i\theta_{a}})-(1-p_{a}e^{i\theta_{a}})\log(1-p_{a}e^{i\theta_{a}})$ (S60)
$-p_{a}e^{-i\theta_{a}}\log(p_{a}e^{-i\theta_{a}})-(1-p_{a}e^{-i\theta_{a}})\log(1-p_{a}e^{-i\theta_{a}})$
$=$ $2p_{a}(\theta_{a}\sin\theta_{a}-\log p_{a}\cos\theta_{a})+2p_{b}(\theta_{b}\sin\theta_{b}-\log p_{b}\cos\theta_{b}),$
@@
where $p_{b},\theta_{b}$ are defined such that $1-p_{a}e^{i\theta_{a}}=p_{b}e^{i\theta_{b}}$. As shown above, $\rm{Im}(S_{\text{pair}})=0$, indicating that the entanglement entropy should be a real value. Any imaginary part of the entanglement entropy from computations must be due to numerical errors.
IV IV. Prospects for measuring negative entanglement through the second Rényi entropy
From the previous section, it was established that the negative entanglement is also manifested generically as negative Rényi entropy. Below, we outline a scheme for measuring the second Rényi entropy, defined in the biorthogonal basis as:
$S_{A}^{(2)}=-\log\text{Tr}\big{[}(\rho_{A})^{2}\big{]},$ (S61)
A known approach [122, 123, 124, 125, 126] for measuring the second Rényi entropy or quantum purity involves the SWAP operator, which exchanges two copies of a quantum state:
[…]
A commonly used corollary [127] is $\text{Tr}(\text{SWAP}\rho_{1}\otimes\rho_{2})=\text{Tr}(\rho_{1}\rho_{2})$, from which the second Rényi entropy can be calculated from the expectation value of the SWAP operator on two-copies of the many body state  [122, 123, 124, 125, 126, 128] as:
$\bra{\psi}\otimes\bra{\psi}\text{SWAP}_{A}\ket{\psi}\otimes\ket{\psi}=\text{Tr}(\text{SWAP}_{A}\rho\otimes\rho)=\text{Tr}(\rho_{A}^{2})\quad=e^{-S_{A}^{(2)}},$ (S63)
[…]
$\big{|}\bra{\Psi^{L}}\otimes\bra{\Psi^{L}}\text{SWAP}_{A}\ket{\Psi^{R}}\otimes\ket{\Psi^{R}}\big{|}$ | $=$ $e^{-\text{Re}[S_{A}^{(2)}]}.$ (S64)
Therefore, to measure the (real part of the) biorthogonal second Rényi entropy, $\text{Re}[S_{A}^{(2)}]$, the most mathematically direct way would be to prepare two copies of the ground state of the Hamiltonian $H$ as |$\ket{\Psi^{R}}\otimes\ket{\Psi^{R}}$, applying the SWAP operator in subregion $A$, and then measuring its overlap with the ground state |$\ket{\Psi^{L}}\otimes\ket{\Psi^{L}}$ of another Hamiltonian $H^{\dagger}$. This approach could potentially be implemented using programmable quantum computers [129, 130, 131, 132]. Post-selection, which has been used in measuring negative conditional entropy [63, 64, 133], will also be useful in simulating the non-Hermiticity [134, 135, 38, 136].
[…]
Thus, to measure $\big{|}\bra{\Psi^{L}}\otimes\bra{\Psi^{L}}\text{SWAP}_{A}\ket{\Psi^{R}}\otimes\ket{\Psi^{R}}\big{|}$ | as given in Eq(S64), one feasible strategy is to prepare a superposition state with known amplitudes $c_{1},c_{2}$ in a physical system:
[…]
and then measure the expectation value $\bra{\Psi}\text{SWAP}_{A}\ket{\Psi}$. The left-right overlap terms (i.e. Eq(S64)) can subsequently be calculated by subtracting the contributions from $\langle\text{SWAP}_{A}\rangle_{RR}$ and $\langle\text{SWAP}_{A}\rangle_{LL}$, which can also be independently measured. Moreover, the OBC spectrum of our model, as specified in Eq. (9), is purely real, leading to identical energies for |$\ket{\Psi^{R}}$ and |$\ket{\Psi^{L}}$ and thus facilitating the preparation of their superposition.
Beyond the approach described above, other potentially feasible ways for observing negative entanglement can involve directly measuring the reduced density matrix to calculate the entanglement entropy through quantum state tomography  [137, 138, 139]. Related quantum simulations through quantum Monte Carlo approaches [125, 140, 141] and ultracold atomic optical lattices [142, 143, 144, 145, 122, 123, 146, 147, 148] can also reveal the associated quantum correlations.

### v4

@@ II.2.3 Scaling of entanglement entropy $S_{A}$ for gapped flat bands
As discussed in subsection B.2, we demonstrated that near $k=0$, the dominant element of $P(k)$ is located in the upper right corner, where $P_{1,2L_{y}}(k)\sim 1/r^{L_{y}}$ with $r=\sqrt{\big{|}\frac{(b_{0}-\cos k)^{B}}{a_{0}}\big{|}}<1$. For $b_{0}>1$, this can still be the case with appropriate choice of $a_{0}$. We have
$\bra{x_{1}}\bar{P}_{1,2L_{y}}\ket{x_{2}}$ $=$ $\frac{1}{L}\sum_{k}e^{ik(x_{1}-x_{2})}P_{1,2L_{y}}(k)$ (73)
$\sim$ $\frac{1}{L}\sum_{k}e^{ik(x_{1}-x_{2})}a_{0}^{L_{y}/2}\frac{1}{(b_{0}-\cos k)^{BL_{y}/2}},$
@@
When $L\ll L_{y}$, (for simplicity, we select $L_{y}$ to be an even number to ensure $BL_{y}/2$ is an integer), the following approximation holds
$(1+x)^{BL_{y}}=\sum_{m}C_{BL_{y}/2}^{m}x^{m}\approx C_{BL_{y}/2}^{M}x^{M}$ (75)
where $m=M$ represents the maximum term in the summation. As verified in Fig. 8(a), $M$ exhibits a linear relation with $L_{y}$, expressed as $M\approx\frac{1}{2}(\gamma L_{y}+\zeta)$ (where $\gamma,\zeta$ are coefficients independent of $L_{y}$). Consequently, we find
$(1+x)^{BL_{y}}$$\approx$ $C_{BL_{y}/2}^{M}x^{M}\sim C_{BL_{y}/2}^{(\gamma L_{y}+\zeta)/2}\left(\frac{\pi^{2}}{2(b_{0}-1)L^{2}}\right)^{(\gamma L_{y}+\zeta)/2}$ (76)
$\sim$ $L^{-(\gamma L_{y}+\zeta)}.$
By substituting into Eqs (74) and (73) and considering only the contribution from $k_{1}=\pi/L$, we have
$\bra{x_{1}}\bar{P}_{1,2L_{y}}\ket{x_{2}}$ $\sim$ $\frac{1}{L}e^{ik_{0}(x_{1}-x_{2})}\frac{a_{0}^{L_{y}/2}}{(b_{0}-1)^{BL_{y}/2}(1+x)^{BL_{y}/2}}$ (77)
$\sim$ $L^{\gamma L_{y}+\zeta-1}.$
As we have demonstrated in the previous subsection B.2, since the elements of $\bar{P}_{1,2L_{y}}$ are proportional to $L^{\gamma L_{y}+\zeta-1}$, we can deduce:
$\text{Tr}(\bar{P}^{2})$$\approx$ $2\text{Tr}(\bar{P}_{1,2L_{y}}\bar{P}_{2L_{y},1})\sim L^{\gamma L_{y}+\zeta-1}$
$\Longrightarrow p_{i}^{1st}$ $\sim$ $L^{\frac{\gamma L_{y}+\zeta-1}{2}},$
$S_{A}$ $\sim$ $-4\times\left(\frac{\gamma L_{y}+\zeta-1}{2}\right)\log L$ (78)
$\sim$ $-(\alpha L_{y}+\beta)\log L,$
where we have redefined $2\gamma$ as $\alpha$ and $2(\zeta-1)$ as $\beta$. In the $L\ll L_{y}$ regime, higher orders of $p_{i}$ are negligible, and we only need to consider the first order. Therefore, we conclude: for $L\ll L_{y}$, $S_{A}\sim-\log L$ (Fig. 8(b)), with the gradient depending linearly on $L_{y}$ (see Fig. 8(c)). Even though our derivation had assumed $L\ll L_{y}$, from the numerics in Fig. 1(e) of the main text, we see that this trend still mostly holds as long as $L<L_{y}$.
When $L\gg L_{y}$, Eq (74) can be approximated as:
$(b_{0}-\cos k_{0})^{BL_{y}/2}$ $=$ $\left(b_{0}-1+\frac{1}{2!}\left(\frac{\pi}{L}\right)^{2}-\frac{1}{4!}\left(\frac{\pi}{L}\right)^{4}+\cdots\right)^{BL_{y}/2}$ (79)
@@
Thus, Eq (73), which represents the elements of the largest matrix block $\bar{P}_{1,2L_{y}}$, can be expressed as
$\bra{x_{1}}\bar{P}_{1,2L_{y}}\ket{x_{2}}\sim\left(\frac{a_{0}}{(b_{0}-1)^{B}}\right)^{L_{y}/2}.$ (80)
Following a similar derivation as Eq (78), we now have
@@
$S_{min}$ $\sim$ $-4\times\frac{L_{y}}{4}\log\left(\frac{a_{0}}{(b_{0}-1)^{B}}\right)$ (81)
$=$ $-L_{y}\log\left(\frac{a_{0}}{(b_{0}-1)^{B}}\right).$
@@
This linear dependence on $L_{y}$ is verified in Fig. 8(d). Empirically, we also see from Fig. 1(e) of the main text that this saturation generally sets in as long as $L$ exceeds $L_{y}$.
Figure 9: Contribution of a complex $p_{i}=|p_{i}|e^{i\theta}$ to the real part of $S_{A}$ for two cases: (a) $|p_{i}|\in[0,1]$ and (b) $|p_{i}|\gg 1$.
Figure 10: Negligible imaginary part for complex $p_{i}$ with Re$(p_{i})>1$ or Re$(p_{i})<0$ in both the 4-band exceptional crossing model (a-c) and the 2-band topological flat-band model (d-f). (a) Replot of the topo, non-Hermitian case from Fig. 1(d). (c) Corresponding scaling of Re($p_{i}$) and Im($p_{i}$) for the complex $p_{i}$ circled in (b). (d) Replot of the $L_{y}=15$ case from Fig. 2(e). (f) Corresponding scaling of Re($p_{i}$) and Im($p_{i}$) for the complex $p_{i}$ circled in (e).
II.3 Further properties of the negative entanglement
II.3.1 Complex $p_{i}$ and the negativity of von Neumann and Rényi entropy
In general non-Hermitian systems, the eigenvalues $p_{i}$ of $\bar{P}$ are complex, making it unintuitive in determining how $p_{i}$ affects the real part of $S_{A}$. In this section, we discuss this issue in detail. Most importantly, we will demonstrate that the negativity in entanglement is not only limited to the von Neumann entropy, but also can be extended to nth-order ($n\geq 2$) Rényi entropies.
[…]
Thus, for the von Neumann entropy, we conclude that for nearly real $p_{i}$, when Re$(p_{i})>1$ or Re$(p_{i})<0$, it contributes negatively to Re$(S_{A})$. This conclusion can also be derived as follows:
$\text{Re}[S_{A}(p_{i})]=\text{Re}[-p_{i}\log(p_{i})-(1-p_{i})\log(1-p_{i}])$ (82)
$=$ $\begin{cases}-p_{i}\log(p_{i})+(p_{i}-1)\log(p_{i}-1)<0,\text{ real }p_{i}>1\\ |p_{i}|\log(|p_{i}|)-(|p_{i}|+1)\log(|p_{i}|+1)<0,\text{ real }p_{i}<0\end{cases}$
Since the above discussion is based on the condition that $p_{i}$ is nearly real, we numerically calculated the values of $p_{i}$ for the two models introduced in our paper. As shown in Figs. 10(c) and (f), we focus primarily on the $p_{i}$ values whose real parts extend beyond the range $[0,1]$ (as highlighted by the circles in Figs. 10 (b) and (e)). We observe that Im$(p_{i})$ remains close to zero, indicating that it is indeed negligible compared to Re$(p_{i})$. As expected, these $p_{i}$ contribute to a negative scaling of Re$(S_{A})$ for both models, as shown in Figs. 10 (a) and (d).
Secondly, we derive how the complex $p_{i}$ contributes to the real part the $n$-th order Rényi entropy. For the case $|p_{i}|\gg 1$, which, as demonstrated in the main article, is induced by the large overlap of different eigenstates, we have:
$p_{i}^{n}+(1-p_{i})^{n}\sim\begin{cases}p_{i}^{n}+O(p_{i}^{n-1}),&\text{for even $n$}\\ p_{i}^{n-1}+O(p_{i}^{n-2}),&\text{for odd $n$}\end{cases}$ (83)
Thus, considering only the dominant term, we approximately have:
$S_{A}^{(n)}(p_{i})$ $=$ $\frac{1}{1-n}\log(p_{i}^{n}+(1-p_{i})^{n})$
$\sim$ $\begin{cases}\frac{n}{1-n}(\log|p_{i}|+i\arg(p_{i})),&\text{for even $n$}\\ -(\log|p_{i}|+i\arg(p_{i})),&\text{for odd $n$}\end{cases},$
$\text{Re}[S_{A}^{(n)}(p_{i})]$ $\sim$ $\begin{cases}\frac{n}{1-n}\log|p_{i}|,&\text{for even $n$}\\ -\log|p_{i}|,&\text{for odd $n$}\end{cases}<0.$ (84)
Hence, unlike the von Neumann entropy, for general complex $p_{i}$ with $|p_{i}|\gg 1$, even if the imaginary part is not negligible, its contribution to the real part of Rényi entropy is always negative.
@@ II.3.1 Complex $p_{i}$ and the negativity of von Neumann and Rényi entropy
II.3.2 PT-symmetry-protected real entanglement entropy $S_{A}$ in the flat band model.
@@
In this section, we demonstrate that both the von Neumann entropy and Rényi entropies are rigorously real for PT-symmetry-protected models, as exemplified by our topological flat band model. This result arises from the fact that the eigenvalues $p_{i}$ of $\bar{P}$ consistently occur in complex conjugate pairs.
@@
First, in the $H_{\text{y-OBC}}(k)$ model of Eq (49), the parameters $a_{0},b_{0},t$ are always real numbers, with $b_{0}\geq 1$. As previously demonstrated, this model can be viewed as a non-Hermitian SSH model that is protected by PT-symmetry, and it features purely real eigenvalues and eigenvectors, $|\psi_{n}^{R}(k)\rangle$ and $\langle\psi_{n}^{L}(k)|$. Therefore, the elements of the projector matrix $P_{\alpha,\beta}(k)=\sum_{n\in\text{occ}}\langle\alpha|\psi_{n}^{R}(k)\rangle\langle\psi_{n}^{L}(k)|\beta\rangle$ are all real, with $\alpha,\beta$ labeling the band indices. Upon expanding these elements into real space, and given that $P_{\alpha,\beta}(k)$ are real, we will obtain
$\langle x_{1}|P_{\alpha,\beta}|x_{2}\rangle=\sum_{k}P_{\alpha,\beta}(k)e^{ik(x_{1}-x_{2})}=\Big{(}\langle x_{2}|P_{\alpha,\beta}|x_{1}\rangle\Big{)}^{*}.$ (85)
@@
$V\bar{P}V^{-1}=\text{conj}(\bar{P}),\text{ with }V=\begin{pmatrix}0&0&0&\cdots&\mathbb{I}\\ 0&0&\cdots&\mathbb{I}&0\\ 0&\cdots&\mathbb{I}&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \mathbb{I}&0&0&\cdots&0\end{pmatrix},$ (86)
@@
indicating that $\bar{P}$ possesses PT-symmetry and its spectrum must consists of complex conjugate pairs $(p,p^{*})$.
For the conjugate pairs $(p,p^{*})=\left(|p|e^{i\theta},|p|e^{-i\theta}\right)$, their contribution to the von Neumann entropy is:
$S_{A}^{\text{pair}}$ $=$ $-|p|e^{i\theta}\log(|p|e^{i\theta})-(1-|p|e^{i\theta})\log(1-|p|e^{i\theta})$
$-|p|e^{-i\theta}\log(|p|e^{-i\theta})-(1-|p|e^{-i\theta})\log(1-|p|e^{-i\theta})$
$2|p|\left(\theta\sin\theta-\log|p|\cos\theta\right)+2|p^{\prime}|(\theta^{\prime}\sin\theta^{\prime}-\log|p^{\prime}|\cos\theta^{\prime}),$
@@
where $|p^{\prime}|,\theta^{\prime}$ are defined such that $1-|p|e^{i\theta}=|p^{\prime}|e^{i\theta^{\prime}}$. As shown above, $\rm{Im}(S_{\text{pair}})=0$, indicating that the von Neumann entropy should be a purely real. And their contribution to the $n-$th order Rényi entropy is:
$\left(S_{A}^{(n)}\right)^{\text{pair}}$ $=$ $\frac{\log\left[p^{n}+(1-p)^{n}\right]}{1-n}+\frac{\log\left[(p^{*})^{n}+(1-p^{*})^{n}\right]}{1-n}$ (88)
$=$ $\frac{1}{1-n}\log(|p^{\prime\prime}|e^{i\theta^{\prime\prime}})+\frac{1}{1-n}\log(|p^{\prime\prime}|e^{-i\theta^{\prime\prime}})$
$=$ $\frac{2}{1-n}\log|p^{\prime\prime}|,$
where we have defined
$p^{n}+(1-p)^{n}=|p^{\prime\prime}|e^{i\theta^{\prime\prime}}.$ (89)
This result shows that $\rm{Im}\left[\left(S_{A}^{(n)}\right)^{\text{pair}}\right]=0$, i.e., the Rényi entropies are purely real. Therefore, any observed imaginary part of the entanglement entropy in numerical computations must be attributed to numerical errors.
II.3.3 Further discussion on the role of $\bar{P}$ eigenvalues.
In many-body free fermion systems, as discussed in this paper, the von Neumann entropy (or Rényi entropy) is a mathematical function of the eigenvalues $p_{i}$ of the truncated projector operator $\bar{P}$. This highlights that the truncated projector operator itself can be regarded as a more fundamental quantity, encapsulating the essential features of the system’s entanglement structure.
As highlighted in previous studies [35], the presence of significant coalescence between the right eigenstates causes the momentum-space projector $P(k)$ to approach singularity. When transformed into real space, this results in the projector operators $P$ and $\bar{P}$ exhibiting slowly decaying, long-range matrix hopping elements. In contrast, in trivial cases with no substantial overlap, $\bar{P}$ is characterized by short-range matrix hopping elements. A striking consequence of these long-range matrix hopping elements is the emergence of unique eigenmodes, referred to as Exceptional Bound States. These modes are distinguished by the real parts of their eigenvalues extending well beyond the typical range of $[0,1]$, ultimately contributing to a negative real component of the entanglement entropy (EE). Thus, the emergence of these modes represents a fundamental and profound consequence of the large overlap.
Hence, even without focusing on the entanglement entropy, the presence of a large eigenstate overlap can be regarded as a physical phenomenon in itself if the $\bar{P}$ matrix elements are taken to be physical hoppings. For instance, by constructing a long-range hopping system i.e. a circuit metamaterial array and designing the Hamiltonian to mimic the topologically protected, non-trivial $\bar{P}$ introduced in this paper, one would observe a novel energy spectrum. This spectrum features special isolated eigenenergies that lie well outside the conventional range of $[0,1]$.
Acknowledgments
We thank Ruizhe Shen and Linhu Li for helpful discussions. This work is supported by the National Research Foundation, Singapore under its QEP2.0 programme (NRF2021-QEP2-02-P09) as well as the Ministry of Education, Singapore (MOE award number: MOE-T2EP50222-0003).